OpenStudy (anonymous):

Given sinx=-4/5 and x is in quadrant 3, What is the value of tan x/2? I have an answer but I'm not sure it's right...

OpenStudy (anonymous):

@hartnn Do you know?

hartnn (hartnn):

first find cos x from, \(\sin^2 x +\cos^2x = 1\) sin x = -4/5 is given...

OpenStudy (anonymous):

I think the answer is -2 because it's in the 3rd quadrant...

hartnn (hartnn):

in 3rd quadrant, tan ratio is positive!

OpenStudy (anonymous):

Oh yea sorry! I meant 2.

hartnn (hartnn):

how u got 2 ?

hartnn (hartnn):

clarification required, is it tan (x/2) or (tan x) /2 ??

OpenStudy (anonymous):

\[\tan(\frac{ x }{ 2 })= \sqrt{\frac{ 1-cosx }{ 1+cosx }}\]

OpenStudy (anonymous):

\[\pm \sqrt{\frac{ 1-(\frac{ -3 }{ 5 }) }{ 1+\frac{ -3 }{ 5 } }}\]

hartnn (hartnn):

yes, correct! from that you will select +2 because tan is +ve in 3rd quadrant...

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

ok, i got that you solved correctly :) welcome ^_^

OpenStudy (anonymous):

It was wrong! It's -2!

hartnn (hartnn):

:O we missed a point!

hartnn (hartnn):

x is in quadrant 3, so, (x/2) will be in quadrant 2 !! and tan is negative in quadrant 2.... silly mistake :(

hartnn (hartnn):

sorry!

OpenStudy (anonymous):

It's ok... :(